# CP Violation of Baryon Decays from $N\pi$ Scatterings

### v1

@@ Abstract
There is a long-standing puzzle that the CP violation (CPV) in the baryon systems has never been well established in experiments, while the CPV of mesons have been observed by decades. In this paper, we propose that the CPV of baryon decays can be generated by the scatterings of a nucleon and a pion into some final states, i.e. $N\pi$ scatterings. Benefited by the fruitful data of $N\pi$ scatterings, we can model-independently analyse the strong phases of $b$-baryon decays using the partial wave amplitudes of $N\pi$ scatterings. This makes a great advantage to predict the CPV of baryon decays. We study the processes of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$, $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$. It is found that the total CPV of the above processes in the invariant mass region of $N\pi$ scatterings are at the order of several percent. More importantly, the regional CPV in the Dalitz plot can reach as large as $20\%$. Therefore, we suggest to measure the CPV of these processes, which has a large possibility to achieve the first observation of CPV in the baryon system.
@@
Introduction.— The matter-antimatter asymmetry in the Universe requires the charge-parity (CP) violation in the microscopic particles Sakharov:1967dj . CP violation (CPV) has been observed in the $K$, $B$ and $D$ mesons sequentially during the past 60 years ParticleDataGroup:2022pth . But it has never been well established in any baryon system, which is a long standing puzzle. In the fact that the visible matter of the universe is dominated by baryons, it is very important to study and search for CPV of baryon systems.
@@
In the Standard Model (SM) of particle physics, baryons do not mix with anti-baryons. CPV of baryons happens only in their decays, which is typically direct CPV. It is well known that the CPV always stems from the interference between tree and penguin diagrams in an individual process, such as $\Lambda_{b}^{0}\to p\pi^{-}$ or $pK^{-}$. However, it has not been observed for the CPV of the above processes even at the level of precision of $\mathcal{O}(1\%)$ LHCb:2018fly . In the multi-body decays, there exist a lot of resonances whose interference can also generate CPV. For example, in the three-body decays of $B^{+}\to\pi^{+}\pi^{+}\pi^{-}$ or $B^{+}\to\pi^{+}K^{+}K^{-}$, the interference between $S$-wave and $P$-wave amplitudes, and the $\pi\pi\leftrightarrow K\bar{K}$ rescattering, contribute to large strong phases and thus giant regional CPV of the Dalitz plot, which could be as large as $(50-60)\%$ LHCb:2019jta ; LHCb:2019xmb ; LHCb:2022fpg ; AlvarengaNogueira:2015wpj ; Cheng:2020ipp . Therefore, it deserves to investigate the multi-body decays of $b$-baryons, such as $\Lambda_{b}^{0}\to pK^{-}\pi^{+}\pi^{-}$ and $p\pi^{-}\pi^{+}\pi^{-}$ which have large data samples at LHCb. If there also exists large regional CPV, it is then possible to observe CPV of baryon decays.
@@
There are a lot of $N^{*}$ excited states in the invariant masses of $p\pi$, $p\pi\pi$ and $\Lambda K$. They could provide fruitful effects of interference and possibly large strong phases, thus possibly large regional CPV. However, there exists a problem meanwhile, that the number of excited $N^{*}$ states are too large. For example, there are more than 15 $N^{*}$ states below 2 GeV. They have different quantum numbers of spin and parity. Their masses and widths are usually of large uncertainties. This makes a significantly difficulty for both experimental partial-wave amplitude analysis and theoretical predictions.
@@
Actually, the excited $N^{*}$ states are usually determined in the experiments of $N\pi$ scatterings. Fortunately, scientists have analyzed the data of $N\pi$ scatterings and provided the partial-wave amplitudes for us. This can be easily obtained from the websites, such as the SAID program SAID . This helps us to model-independently calculate the strong phases of $\Lambda_{b}$ decays, regardless of the detailed information of the excited states. It naturally solves the problem of too many resonances and large uncertainties from the excited $N^{*}$ states. The strong phases are usually of the most difficulty and the largest uncertainty in the predictions of CPV in theory. But using the data of scatterings can significantly improve the theoretical precision on the predictions of CPV. For example, the $\pi\pi\leftrightarrow K\bar{K}$ rescattering data leads to a very accurate description of CPV in $B^{\pm}\to K^{\pm}\pi^{+}\pi^{-}$ and $K^{\pm}K^{+}K^{-}$ Garrote:2022uub , and the enhancement of CPV of $D^{0}\to K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Bediaga:2022sxw . Analogously, the partial-wave amplitudes obtained in the data of $N\pi\to p\pi$ or $p\pi\pi$ scatterings would be helpful to achieve a model-independent determination of strong phases, thus improve the precision of predictions on CPV of $\Lambda_{b}$ decays.
@@
In this Letter, we propose that there might be large regional CPV in the decays of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$, $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$. The mechanism of CPV in these processes is depicted in Fig. 1. In the short-distance weak interaction, $\Lambda_{b}^{0}$ transited into $N\pi$ with an emitted $\pi^{-}$ or $K^{-}$. Then $N\pi$ scatters into $p\pi^{+}\pi^{-}$ or $p\pi^{0}$ via a long-distance strong interaction. It can be model-independently obtained for the partial-wave amplitudes of scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$ from the data analysis by the SAID program SAID . This is the most important feature of this proposal. The interference effect between the partial-wave amplitudes are different for the tree and penguin operators, which results in the strong phase difference between the tree and penguin amplitudes of $\Lambda_{b}$ decays. The strong phases from $N\pi$ scatterings are usually large. So it is possible to generate large regional CPV.
@@
This is the first proposal for the CPV mechanism of baryon decays stemming from $N\pi$ scatterings. On one hand, the strong phases can be model-independently obtained, so that the theoretical predictions are more reliable. It realizes the remarkable dynamical predictions of CPV in $\Lambda_{b}$ decays with long-distance contributions. On the other hand, we can do the angular analysis since the data of $N\pi$ scatterings provides very comprehensive partial-wave amplitudes. The interference between different partial-wave amplitudes can be revealed in the angular distributions. From a detailed numerical analysis, we predict both the global and regional CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$, in the mass regions of scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$. It is found that the global CPV of these processes are several percent, while the regional CPV could reach as large as $(20-30)\%$. It could be even larger for the CPV of some moments of $\cos\theta$ in the angular distributions. Some suggestions are then given for the experimental measurements, which are of large possibility for the first observation of baryon CPV.
CPV mechanism from $N\pi$ scatterings.— The CPV mechanism of baryon decays from the $N\pi$ scatterings is depicted in Fig.1.
Figure 1: Illustration of the CPV mechanism driven by $N\pi$ scatterings. The weak phases come from the tree and penguin operators, while the strong phases stem from the $N\pi$ scatterings. The processes are $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$.
CPV comes from the interference between the figures of Figs. 1 (a) and (b). The weak phases are from tree and penguin operators, while the strong phases stem from the $N\pi$ scatterings except for the effective Wilson coefficients.
@@
The decay amplitudes of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ undergo two steps, $\mathcal{A}=\mathcal{S}^{1/2}\mathcal{A}_{0}$, given by the decay mechanism of final-state interaction of weak decays of heavy hadrons Suzuki:1999uc ; Cheng:2004ru ; Chua:2007cm . $\mathcal{A}_{0}$ is the short-distance weak decays of $\Lambda_{b}^{0}\to(N\pi)h^{-}$, while $\mathcal{S}^{1/2}$ represents the long-distance scatterings of $N\pi\to p\pi^{+}\pi^{-}$ or $p\pi^{0}$. The CPV mechanism of $N\pi$ scatterings is similar to the $\pi\pi\to K\bar{K}$ rescattering in $B$-meson three-body decays Bediaga:2013ela ; AlvarengaNogueira:2015wpj ; Cheng:2016shb ; Garrote:2022uub . In $B$-meson decays, the rescattering of $\pi\pi\to K\bar{K}$ happens between two channels, like $B^{\pm}\to\pi^{\pm}\pi^{+}\pi^{-}$ and $\pi^{\pm}K^{+}K^{-}$. The $N\pi$ scatterings in $\Lambda_{b}$ decays contribute to one individual process but with different partial waves.
Firstly, $\Lambda_{b}$ decays via a weak interaction at a very short time, with $b\to u\bar{u}q$, $q=d/s$. This is a short-distance dynamics due to the large mass scale of $b$ quark and the weak interaction. In the factorization hypothesis, it can be naively separated into $\Lambda_{b}\to N\pi$ transition form factors and the decay constants of emitted $h^{-}=\pi^{-}/K^{-}$ in both the tree and penguin diagrams. The full expression of the transition form factors of $\Lambda_{b}\to N\pi$ is long and not necessary to be given at the current stage. We will only consider the $\gamma_{\mu}(\gamma_{5})$ terms for convenience. In order to connect with the $N\pi$ scatterings in the second step, the $N\pi$ systems are taken as a whole with different spin-parity spinors. The matrix element of transition from $\Lambda_{b}$ into $N\pi$ is
$\langle N\pi|(\bar{u}b)_{\rm V-A}|\Lambda_{b}\rangle$
$=\bar{u}_{N\pi}^{1/2^{+}}(f_{1}^{1/2^{+}}\gamma_{\mu}-g_{1}^{1/2^{+}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
$+\bar{u}_{N\pi}^{1/2^{-}}(-f_{1}^{1/2^{-}}\gamma_{\mu}+g_{1}^{1/2^{-}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
$+\bar{u}_{N\pi}^{3/2^{+}}(f_{1}^{3/2^{+}}\gamma_{\mu}-g_{1}^{3/2^{+}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
$+\bar{u}_{N\pi}^{3/2^{-}}(-f_{1}^{3/2^{-}}\gamma_{\mu}+g_{1}^{3/2^{-}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
++$\cdots$,+ ⋯ , (1)
where the form factors and the spinors are labelled by the spin-parity quantum numbers of the $N\pi$ systems.
@@
The $N\pi$ system produced via short-distance interaction undergoes a long-distance scattering process, i.e. the right-handed parts of Figs. 1 (a) and (b). We consider the scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$. In practice, $p\pi^{+}\pi^{-}$ is dominated by $\Delta^{++}\pi^{-}$. If necessary, it could be cut for the mass of region of $p\pi^{+}$ to be around $\Delta^{++}(1232)$ whose next excited state $\Delta(1600)$ is very far from it. So we will discuss $\Delta^{++}\pi^{-}$ instead of $p\pi^{+}\pi^{-}$ for the convenience. There are fruitful experimental data of the scatterings of $N\pi\to N\pi$ and $\Delta\pi$. Scientists have provided the real and imaginary parts of the partial-wave amplitudes of these scatterings, as an opne source in the website, such as the SAID program SAID . The details of the partial-wave amplitudes can be found in the website. This is most important point of this work. The long-distance contributions are usually very difficult to be calculated due to the non-perturbative property. It would significantly reduce the theoretical uncertainty if directly use the experimental data of scatterings. Besides, there are a lot of excited $N^{*}$ states whose masses and widths are of large uncertainties. The using of data of $N\pi$ scatterings would avoid the complex and unclear information of the resonant states. The real and imaginary parts of the partial-wave amplitudes contribute the strong phases which are necessary for CPV.
@@
Besides, the isospin of the $N\pi$ system can only be 1/2, since the initial state $\Lambda_{b}$ is isospin singlet, and the weak transition of $b\to u$ changes the isospin by $\Delta I=1/2$. Therefore, only the partial-wave amplitudes with isospin 1/2 contribute to the $N\pi$ scatterings in our case. We will use the same symbols as the SAID program to represent the partial-wave amplitudes, $L_{2I,2J}$. $L$ is the angular momentum between $N$ and $\pi$, while $I$ and $J$ are the isospin and total spin of the system. For example, $D_{13}$ represents a partial-wave amplitude with $L=2$, $I=1/2$ and $J=3/2$. In fact, the spin-parity quantum states and the partial-wave amplitudes are one-to-one corresponding to each other, such as $1/2^{+}:P_{11}$, $1/2^{-}:S_{11}$, $3/2^{+}:P_{13}$, $3/2^{-}:D_{13}$, etc.
@@
where $\lambda_{i}=V_{ib}V_{iq}^{*}$ with $i=u,t$ and $q=d,s$, $f_{h}$ is the decay constant of $h=\pi$ or $K$, $a_{46\pm}=a_{4}\pm R_{h}a_{6}$, $R_{h}=2m_{h}^{2}/(m_{b}(m_{u}+m_{q}))$, and $a_{1,4,6}$ are the effective Wilson coefficients in the weak Hamiltonian Buchalla:1995vs ; Ali:1998eb , $m_{\pm}=m_{\Lambda_{b}}\pm m_{N\pi}$.

### v3

# CP Violation of Baryon Decays with $N\pi$ Rescatterings

@@ Abstract
There is a long-standing puzzle that the CP violation (CPV) in the baryon systems has never been well established in experiments, while the CPV of mesons have been observed by decades. In this paper, we propose that the CPV of baryon decays can be generated with the rescatterings of a nucleon and a pion into some final states, i.e. $N\pi\to N\pi$ or $N\pi\pi$. Benefited by the fruitful data of $N\pi$ scatterings, we can model-independently analyse the strong phases of $b$-baryon decays using the partial wave amplitudes of $N\pi$ scatterings. Avoiding the most difficult problem of non-perturbative dynamics, it makes a great advantage to predict the CPV of baryon decays with a relatively reliable understanding of the decay dynamics. We study the processes of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$. It is found that the global CPV of the above processes in the invariant mass regions of $N\pi$ scatterings are at the order of several percent. More importantly, the local CPV in some regions of the Dalitz plots can reach the order of $10\%$, or be even larger. Considering the predicted results and the experimental data samples, we strong suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$, which has a large possibility to achieve the first observation of CPV in the baryon system.
I Introduction
@@
The matter-antimatter asymmetry in the Universe requires the violation of the charge-parity (CP) symmetry in the microscopic particles and anti-particles Sakharov:1967dj . CP violation (CPV) has been observed in the $K$, $B$ and $D$ mesons sequentially during the past 60 years ParticleDataGroup:2022pth . But it has never been well established in any baryon system, which is a long standing puzzle. In the fact that the visible matter of the universe is dominated by baryons, it is very important to study and search for CPV of baryon systems.
@@
In the Standard Model (SM) of particle physics, baryons do not mix with anti-baryons. CPV of baryons happens only in their decays, which is typically direct CPV. It is well known that CPV always stems from the interference between tree and penguin diagrams in an individual process, such as $\Lambda_{b}^{0}\to p\pi^{-}$ or $pK^{-}$. However, it has not been observed for the CPV of the above processes even at the level of precision of $\mathcal{O}(1\%)$ LHCb:2018fly . In the multi-body decays, there exist a lot of resonances whose interference can also generate CPV. For example, in the three-body decay of $B^{+}\to\pi^{+}\pi^{+}\pi^{-}$, the interference between $f_{0}(500)$ and $\rho(770)$, or say the interference between the $S$-wave and $P$-wave amplitudes, contributes to the local CP violations as large as $(50-60)\%$ in some regions of the Dalitz plot LHCb:2019jta ; LHCb:2022fpg . Therefore, it deserves to investigate the multi-body decays of $b$-baryons.
@@
A lot of baryon resonances can contribute to the multi-body decays of $b$-baryons. For instance, the excited states of nucleons $N^{*}$ with different spins and parities could provide fruitful effects of interference in the invariant masses of $p\pi$ or $p\pi\pi$. They might contribute to large strong phases, and thus possibly large regional CPV. However, there exists a problem meanwhile, that the number of excited $N^{*}$ states are too large. For example, it is recorded for more than 15 $N^{*}$ states in the mass region from the first excited state $N(1440)$ to 2 GeV in the Particle Data Group (PDG) ParticleDataGroup:2022pth . Their widths are usually larger than 100 MeV, making it very difficult to distinguish so many states within the invariant mass range of 600 MeV. Besides, the masses and widths of these states are usually of large uncertainties in PDG. This makes a significant difficulty for both experimental partial-wave amplitude analysis and theoretical predictions.
@@
Actually, the excited $N^{*}$ states are usually determined in the experiments of $N\pi$ scatterings. Fortunately, scientists have analyzed the data of $N\pi$ scatterings and provided the partial-wave amplitudes. This can be easily obtained from the websites, such as the SAID program SAID . This helps us model-independently calculate the strong phases of $\Lambda_{b}$ decays, regardless of the detailed information of the resonant states in the invariant masses of $p\pi$ or $p\pi\pi$. It naturally solves the problem of too many resonances with large widths and large uncertainties from the excited $N^{*}$ states. The strong phases are usually of the most difficulty and of the largest uncertainty in the predictions of CPV in theory. But using the data of scatterings can significantly improve the theoretical precision on the predictions of CPV. For example, the $\pi\pi\leftrightarrow K\bar{K}$ rescattering data leads to a very accurate description of CPV in $B^{\pm}\to K^{\pm}\pi^{+}\pi^{-}$ and $K^{\pm}K^{+}K^{-}$ AlvarengaNogueira:2015wpj ; Cheng:2020ipp ; Garrote:2022uub , and the enhancement of CPV of $D^{0}\to K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Bediaga:2022sxw . Analogously, the partial-wave amplitudes obtained in the data of $N\pi\to p\pi$ or $p\pi\pi$ scatterings would be helpful to achieve a model-independent determination of strong phases, thus improve the precision of predictions on CPV of $\Lambda_{b}$ decays.
@@
In this work, we propose that there might be large local CPV in the decays of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$, in the mass region of $p\pi^{+}\pi^{-}$ or $p\pi^{0}$ re-scattered from $N\pi$. The mechanism of CPV in these processes is depicted in Fig. 1. In the short-distance weak interaction, $\Lambda_{b}^{0}$ transited into $N\pi$ with an emitted $\pi^{-}$ or $K^{-}$. Then $N\pi$ scatters into $p\pi^{+}\pi^{-}$ or $p\pi^{0}$ via a long-distance strong interaction. It can be model-independently obtained for the partial-wave amplitudes of scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$ from the data analysis by the SAID program SAID . This is the most important feature of this proposal. The interference effect between the partial-wave amplitudes are different for the tree and penguin operators, which results in the strong phase differences between the tree and penguin amplitudes of $\Lambda_{b}$ decays. The strong phases from $N\pi$ scatterings are usually large. So it is possible to generate large regional CPV.
@@
This is the first proposal for the CPV mechanism of baryon decays stemming from $N\pi$ scatterings. On one hand, the strong phases are data driven and can thus be model-independently obtained, so that the theoretical predictions are more reliable. It realizes the remarkable dynamical predictions of CPV in $\Lambda_{b}$ decays with long-distance contributions which are usually very difficult in theory Lu:2009cm ; Hsiao:2014mua ; Zhu:2018jet . On the other hand, we can do the angular analysis since the data of $N\pi$ scatterings provides very comprehensive partial-wave amplitudes. The interference between different partial-wave amplitudes can be revealed in the angular distributions. From a detailed numerical analysis, we predict both the global and regional CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$, in the mass regions of scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$. It is found that the global CPV of these processes are several percent, while the regional CPV could reach as large as $(20-30)\%$. It could be even larger for the CPV of some moments of $\cos\theta$ in the angular distributions. Some suggestions are then given for the experimental measurements, which are of large possibility for the first observation of baryon CPV.
II CPV mechanism with $N\pi$ scatterings
The CPV mechanism of baryon decays via the $N\pi$ scatterings is depicted in Fig.1. Take the decays of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ as an example. The decay amplitudes of these processes undergo two steps,
$\mathcal{A}=\mathcal{S}^{1/2}\mathcal{A}_{0},$ (1)
@@ II CPV mechanism with $N\pi$ scatterings
given by the decay mechanism of final-state interaction of weak decays of heavy hadrons Suzuki:1999uc ; Cheng:2004ru ; Chua:2007cm . $\mathcal{A}_{0}$ is the short-distance weak decays of $\Lambda_{b}^{0}\to(N\pi)h^{-}$, while $\mathcal{S}^{1/2}$ represents the long-distance final-state re-scatterings of $N\pi\to p\pi^{+}\pi^{-}$ or $p\pi^{0}$.
Figure 1: Illustration of the CPV mechanism driven by $N\pi$ re-scatterings. The processes undergo two steps, the short-distance weak interaction of $b$ decays firstly, and then long-distance strong interactions of final-state re-scatterings of $N\pi\to N\pi$ or $N\pi\pi$. The weak phases come from the tree and penguin diagrams, while the strong phases stem from the partial-wave amplitudes of $N\pi$ scatterings which can be model-independently obtained from the data of the SAID program. For example, the re-scatterings of $N\pi\to p\pi^{0}$ and $p\pi^{+}\pi^{-}$ contribute to the processes of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$, respectively.
CPV comes from the interference between the diagrams of 1 (a) and (b). The weak phases are from the tree and penguin processes, while the strong phases stem from the $N\pi$ scatterings associated with the effective Wilson coefficients.
Firstly, $\Lambda_{b}$ decays into $(N\pi)h^{-}$ via a weak interaction of $b\to u\bar{u}d(s)$ at a very short time, i.e. the left-handed parts of Figs. 1 (a) and (b). This is a short-distance dynamics due to the large mass scale of $b$ quark and the weak interaction. The $N\pi$ system can have all possible quantum numbers of spin and parity. The decay amplitudes of $\Lambda_{b}\to(N\pi)h^{-}$ with short-distance dynamics are expressed as
$\mathcal{A}_{0}=$ $\hskip 8.5359pt\bar{u}_{N\pi,1/2^{+}}(A+B\gamma_{5})u_{\Lambda_{b}}$
$+\bar{u}_{N\pi,1/2^{-}}(\tilde{A}+\tilde{B}\gamma_{5})u_{\Lambda_{b}}$
$+q_{\mu}\bar{u}_{N\pi,3/2^{+}}^{\mu}(C+D\gamma_{5})u_{\Lambda_{b}}$
$+q_{\mu}\bar{u}_{N\pi,3/2^{-}}^{\mu}(\tilde{C}+\tilde{D}\gamma_{5})u_{\Lambda_{b}}$
++$\cdots$.+ ⋯ . (2)
The spinors of final states correspond to the $N\pi$ system with different spins and parities, instead of that of the nucleon $N$. The $N\pi$ system is also a fermion which can be described by spinors. Such expression is more convenient to match with the $N\pi$ rescatterings in the following discussions. The spinors of spin-1/2 system is simple, while those of spin-3/2 are Rarita-Schwinger spinors with Lorentz indices. $q_{\mu}$ is the four momentum of $h^{-}$. The detailed expressions between the processes with the same spin but different parities are similar to each other, thus are distinguished only by tilde in the decay amplitudes of $A$, $B$ and $C$, $D$. It is noted that the above expressions are stringent, with only two terms with or without $\gamma_{5}$ in the decay amplitude for each process, due to the simplicity of the emitted pseudoscalar meson $h=\pi$ or $K$. The $N\pi$ system with higher spins are all given in a similar way, but not explicitly expressed here.
A detailed dynamics of the decay amplitudes could be calculated under the factorization hypothesis. It is well known that the factorization hypothesis works well for the external $W$-emission diagram, and this diagram is dominated compared to the other topological diagrams in the non-leptonic weak decays of heavy-flavor hadrons. This can be seen in the decays of $B$-mesons Ali:1998eb , $b$-baryonsbaryonPQCD , $D$-mesonsLi:2012cfa and charmed baryonsFSI-charm . In the factorization hypothesis, the decay amplitudes can be naively separated into the transition form factors of $\Lambda_{b}\to N\pi$ and the decay constants of emitted $h^{-}$ in both the tree and penguin diagrams, multiplied by the corresponding CKM matrix elements and effective Wilson coefficients. The full expressions of the $\Lambda_{b}\to N\pi$ form factors are long and not necessary to be explicitly given at the current stage. The form factors with the spin-parity states of $1/2^{\pm}$, $3/2^{\pm}$ and $5/2^{\pm}$ can be refereed to Mott:2011cx ; Gutsche:2017wag ; Meinel:2021mdj ; Huang:2022lfr .
@@
Secondly, the $N\pi$ system produced via short-distance weak interaction undergoes a long-distance re-scattering process, i.e. the right-handed parts of Figs. 1 (a) and (b). We only consider the re-scatterings of $N\pi\to p\pi^{0}$ and $p\pi^{+}\pi^{-}$ in this work to illustrate this novel CPV mechanism. There are fruitful experimental data of the scatterings of $N\pi\to N\pi$ and $N\pi\pi$. Scientists have provided the real and imaginary parts of the partial-wave amplitudes of these scatterings, as an open source in the website such as the SAID program SAID . The details of the partial-wave amplitudes can be found in the website, and shown with some examples in the Appendix A. This is the most important point of this work. The long-distance contributions are usually very difficult to be calculated due to the non-perturbative property. However, the use of the experimental data of scatterings could significantly reduce the theoretical uncertainties in the calculations of decay amplitudes. Besides, there are a lot of excited $N^{*}$ states whose masses and widths are of large uncertainties. The data of $N\pi$ scatterings would avoid the complicated and unclear information of these resonances. The real and imaginary parts of the partial-wave amplitudes contribute to the strong phases which are necessary for CPV.
In practice, $p\pi^{+}\pi^{-}$ is dominated by $\Delta^{++}\pi^{-}$ LHCb:2019jyj . So we will discuss $\Delta^{++}\pi^{-}$ instead of $p\pi^{+}\pi^{-}$ for the convenience. It is found by the LHCb collaboration in LHCb:2019jyj that the signal yields of $\Lambda^{0}_{b}\to\Delta^{++}\pi^{-}\pi^{-}$ and $\Delta^{++}\pi^{-}K^{-}$ are significantly large. There are some other scattering processes, such as $N\pi\to N\rho\to p\pi\pi$, whose partial-wave amplitudes are also provided by the SAID program. However, we would like to leave the comprehensive analysis of $N\pi\to p\pi^{+}\pi^{-}$ in a detailed study in the near future. The purpose of this work is not to do a very comprehensive numerical predictions, but focus on the illustration of this new CPV mechanism via $N\pi$ scatterings. Therefore, it is good enough to use an approximation of the $\Delta^{++}\pi^{-}$ dominance in $p\pi^{+}\pi^{-}$ as a starting point.
@@
Besides, the isospin of the $N\pi$ system can only be 1/2, since the initial state $\Lambda_{b}$ is isospin singlet, and the weak transition of $b\to u$ changes the isospin by $\Delta I=1/2$. Therefore, only the partial-wave amplitudes with isospin 1/2 contribute to the $N\pi$ scatterings in our case. We will use the same symbols as the SAID program to represent the partial-wave amplitudes, $L_{2I,2J}$. $L$ is the angular momentum between $N$ and $\pi$, while $I$ and $J$ are the isospin and total spin of the system. For example, $D_{13}$ represents a partial-wave amplitude with $L=2$, $I=1/2$ and $J=3/2$. In fact, the spin-parity quantum states and the partial-wave amplitudes are one-to-one corresponding to each other, such as $1/2^{+}:P_{11}$, $1/2^{-}:S_{11}$, $3/2^{+}:P_{13}$, $3/2^{-}:D_{13}$, etc. Note that the partial-wave rescattering amplitudes depend on the invariant masses of $N\pi$. The details of the partial-wave amplitudes of $N\pi$ scatterings are shown in Appendix.A for the convenience. It can be found that the real and imaginary parts of the scattering amplitudes are of high precision, and thereby the strong phases can be precisely obtained. According to the final-state rescattering formula of Eq. (1), the total decay amplitudes are expressed as
$\mathcal{A}=$ $\hskip 8.5359pt\bar{u}_{N\pi,1/2^{+}}(A+B\gamma_{5})u_{\Lambda_{b}}\;P_{11}$
$+\bar{u}_{N\pi,1/2^{-}}(\tilde{A}+\tilde{B}\gamma_{5})u_{\Lambda_{b}}\;S_{11}$
$+q_{\mu}\bar{u}_{N\pi,3/2^{+}}^{\mu}(C+D\gamma_{5})u_{\Lambda_{b}}\;P_{13}$
$+q_{\mu}\bar{u}_{N\pi,3/2^{-}}^{\mu}(\tilde{C}+\tilde{D}\gamma_{5})u_{\Lambda_{b}}\;D_{13}$
++$\cdots$.+ ⋯ . (3)
Thirdly, we illustrate the CPV mechanism with $N\pi$ scatterings. Since the transition form factors of $\Lambda_{b}\to N\pi$ have not been calculated up to now, we will only consider the terms with $\gamma_{\mu}$ and $\gamma_{\mu}\gamma_{5}$. For example, the transition form factors of $\Lambda_{b}$ into $J^{P}=1/2^{\pm}$ baryons are dominated by the terms of $\bar{u}_{1/2^{+}}(f_{1}^{1/2^{+}}\gamma_{\mu}+g_{1}^{1/2^{+}}\gamma_{\mu}\gamma_{5})u_{\Lambda_{b}}$ and $-\bar{u}_{1/2^{-}}(f_{1}^{1/2^{-}}\gamma_{\mu}+g_{1}^{1/2^{-}}\gamma_{\mu}\gamma_{5})u_{\Lambda_{b}}$ Han:2022srw ; Huang:2022lfr , respectively, for the V-A current. The minus sign for the $1/2^{-}$ process is from the negative parity, with the definition of the form factors $f_{1}^{1/2^{-}}$ and $g_{1}^{1/2^{-}}$ as positive values. The form factors with higher-spin $N\pi$ states have similar properties.
Except for neglecting the contributions from other terms of form factors, we also ignore the sub-leading diagrams such as non-factorizable diagrams in which $h^{-}$ could also re-scatter with $N$ or $\pi$. The purpose of this paper is to find a mechanism which can drive the CPV of baryon decays in the low mass region of $p\pi$ and $p\pi\pi$, especially for generating large regional CPV. Therefore, we investigate the main mechanism and consider only the dominant contributions, using the above approximations without big problems and leaving the ignored effects in the further discussions in the future.
@@
where $\lambda_{i}=V_{ib}V_{iq}^{*}$ with $i=u,t$ and $q=d,s$, $f_{h}$ is the decay constant of $h=\pi$ or $K$, $a_{46\pm}=a_{4}\pm R_{h}a_{6}$, $R_{h}=2m_{h}^{2}/(m_{b}(m_{u}+m_{q}))$, and $a_{1,4,6}$ are the effective Wilson coefficients in the weak Hamiltonian Ali:1998eb , $m_{\pm}=m_{\Lambda_{b}}\pm m_{N\pi}$.